### v2

@@ Acknowledgements.
This work benefited from support through Project Nos. GR 967/7-1 and GI 1405/1-1 of the Deutsche Forschungsgesellschaft (DFG). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Ján Šimkanin is grateful for support from the Alexander-von-Humboldt Stiftung (CZE 1079936).References
@@
- Stefani et al. [2019] F. Stefani, A. Gailitis, G. Gerbeth, A. Giesecke, T. Gundrum, G. Rüdiger, M. Seilmayer, and T. Vogt, “The dresdyn project: liquid metal experiments on dynamo action and magnetorotational instability,” Geophys. Astrophys. Fluid Dyn. 113, 51 – 70 (2019).
@@
- Giesecke et al. [2024] A. Giesecke, T. Vogt, F. Pizzi, V. Kumar, F. Garcia-Gonzalez, T. Gundrum, and F. Stefani, “The global flow state in a precessing cylinder.” J. Fluid Mech. XXX, in press (2024).
@@
- Gans [1971] R. F. Gans, “On hydromagnetic precession in a cylinder,” J. Fluids Mech. 45, 111–130 (1971).
@@
- Blackburn and Sherwin [2004] H. M. Blackburn and S. J. Sherwin, “Formulation of a Galerkin spectral element-Fourier method for three-dimensional incompressible flows in cylindrical geometries,” J. Comp. Phys. 197, 759–778 (2004).
- Blackburn et al. [2019] H. M. Blackburn, D. Lee, T. Albrecht, and J. Singh, “Semtex: a spectral element–Fourier solver for the incompressible Navier–Stokes equations in cylindrical or Cartesian coordinates,” Comput. Phys. Cmm. 245, 106804–1–13 (2019).
- Giesecke et al. [2010] A. Giesecke, C. Nore, F. Stefani, G. Gerbeth, J. Léorat, F. Luddens, and J.-L. Guermondz, “Electromagnetic induction in non-uniform domains,” Geophys. Astrophys. Fluid Dyn. 104, 505 – 529 (2010).
- Giesecke [2024a] A. Giesecke, “supplemental material,” www.giesecke.net (2024a).
- Giesecke [2024b] A. Giesecke, “supplemental material,” www.giesecke.net (2024b).
- Rüdiger, Elstner, and Ossendrijver [2003] G. Rüdiger, D. Elstner, and M. Ossendrijver, “Do spherical -dynamos oscillate?” Astron. Astrophys. 406, 15–21 (2003).

### v3

@@ Acknowledgements.
This work benefited from support through Project Nos. GR 967/7-1 and GI 1405/1-1 of the Deutsche Forschungsgesellschaft (DFG). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Ján Šimkanin is grateful for support from the Alexander-von-Humboldt Stiftung (CZE 1079936) and for funding from the Helmholtz Association in frame of the AI project GEOMAGFOR (ZT-I-PF-5-200).References
@@
- Stefani et al. [2015] F. Stefani, T. Albrecht, G. Gerbeth, A. Giesecke, T. Gundrum, J. Herault, C. Nore, and C. Steglich, “Towards a precession driven dynamo experiment,” Magnetohydrodynamics 51, 275–284 (2015).
@@
- Giesecke et al. [2024] A. Giesecke, T. Vogt, F. Pizzi, V. Kumar, F. Garcia-Gonzalez, T. Gundrum, and F. Stefani, “The global flow state in a precessing cylinder.” J. Fluid Mech. 998, A30 (2024).
- Kobine [1996] J. J. Kobine, “Azimuthal flow associated with inertial wave resonance in a precessing cylinder,” J. Fluid Mech. 319, 387 – 406 (1996).
- Manasseh [1992] R. Manasseh, “Breakdown regimes of inertia waves in a precessing cylinder,” J. Fluids Mech. 243, 261–296 (1992).
- Manasseh [1994] R. Manasseh, “Distortions of inertia waves in a rotating fluid cylinder forced near its fundamental mode resonance,” J. Fluids Mech. 265, 345–370 (1994).
- Manasseh [1996] R. Manasseh, “Nonlinear behaviour of contained inertia waves,” J. Fluids Mech. 315, 151–173 (1996).
- Noir et al. [2001] J. Noir, D. Brito, K. Aldridge, and P. Cardin, “Experimental evidence of inertial waves in a precessing spheroidal cavity,” Geophys. Res. Lett. 28, 3785–3788 (2001).
- Noir et al. [2003] J. Noir, P. Cardin, D. Jault, and J.-P. Masson, “Experimental evidence of non-linear resonance effects between retrograde precession and the tilt-over mode within a spheroid,” Geophys. J. Int. 154, 407–416 (2003).
- Mouhali et al. [2012] W. Mouhali, T. Lehner, J. Léorat, and R. Vitry, “Evidence for a cyclonic regime in a precessing cylindrical container,” Exp. Fluids 53, 1693–1700 (2012).
[…]
- Le Bars et al. [2022] M. Le Bars, A. Barik, F. Burmann, D. P. Lathrop, J. Noir, N. Schaeffer, and S. A. Triana, “Fluid dynamics experiments for planetary interiors,” Surveys in Geophysics 43, 229 – 261 (2022).
- Burmann, Kira, and Noir [2024] F. Burmann, L. Kira, and J. Noir, “Precessing non-axisymmetric ellipsoids: bi-stability and fluid instabilities,” Journal of Fluid Mechanics 997 (2024), 10.1017/jfm.2024.774.
- Christensen and Aubert [2006] U. R. Christensen and J. Aubert, “Scaling properties of convection-driven dynamos in rotating spherical shells and application to planetary magnetic fields,” Geophysical Journal International 166, 97–114 (2006).
- King et al. [2010] E. M. King, K. M. Soderlund, U. R. Christensen, J. Wicht, and J. M. Aurnou, “Convective heat transfer in planetary dynamo models,” Geochemistry, Geophysics, Geosystems 11 (2010), https://doi.org/10.1029/2010GC003053.
@@
- Cappanera et al. [2016] L. Cappanera, J.-L. Guermond, J. Léorat, and C. Nore, “Two spinning ways for precession dynamo,” Phys. Rev. E 93, 043113 (2016).
@@
- Pizzi et al. [2022] F. Pizzi, A. Giesecke, J. Šimkanin, V. Kumar, T. Gundrum, and F. Stefani, “Numerical and theoretical framework for the dresdyn precession dynamo experiment,” Magnetohydrodynamics 58, 445 – 453 (2022).
- Kumar et al. [2023] V. Kumar, F. Pizzi, A. Giesecke, J. Šimkanin, T. Gundrum, M. Ratajczak, and F. Stefani, “The effect of nutation angle on the flow inside a precessing cylinder and its dynamo action,” Phys. Fluids 35 (2023), 10.1063/5.0134562.
- Guermond et al. [2009] J.-L. Guermond, R. Laguerre, J. Léorat, and C. Nore, “Nonlinear magnetohydrodynamics in axisymmetric heterogeneous domains using a fourier/finite element technique and an interior penalty method,” Journal of Computational Physics 228, 2739 – 2757 (2009).
- Giesecke et al. [2010] A. Giesecke, C. Nore, F. Stefani, G. Gerbeth, J. Léorat, F. Luddens, and J.-L. Guermond, “Electromagnetic induction in non-uniform domains,” Geophys. Astrophys. Fluid Dyn. 104, 505 – 529 (2010).
- Rädler et al. [2002] K.-H. Rädler, M. Rheinhardt, E. Apstein, and H. Fuchs, “On the mean-field theory of the karlsruhe dynamo experiment i. kinematic theory,” Magnetohydrodynamics 38, 41 – 71 (2002).
@@
- Dudley and James [1989] M. L. Dudley and R. W. James, “Time-dependent kinematic dynamos with stationary flows,” Proceedings of the Royal Society of London Series A 425, 407–429 (1989).
- Giesecke [2024a] A. Giesecke, “supplemental material,” will be inserted by publisher (2024a).
@@
- Giesecke [2024b] A. Giesecke, “supplemental material,” will be inserted by publisher (2024b).
@@
[…]
- Reddy, Favier, and Le Bars [2018] K. S. Reddy, B. Favier, and M. Le Bars, “Turbulent kinematic dynamos in ellipsoids driven by mechanical forcing,” Geophysical Research Letters 45, 1741 – 1750 (2018).
@@
- Avalos-Zuniga, Plunian, and Gailitis [2003] R. Avalos-Zuniga, F. Plunian, and A. Gailitis, “Influence of electromagnetic boundary conditions onto the onset of dynamo action in laboratory experiments,” Phys. Rev. E 68, 0663071 (2003).
- Avalos-Zuñiga and Plunian [2005] R. Avalos-Zuñiga and F. Plunian, “Influence of inner and outer walls electromagnetic properties on the onset of a stationary dynamo,” European Physical Journal B 47, 127–135 (2005).